Find the number of sides of a regular polygon if the measure of each interior angle is 17 times the measure of each exterior angle.
Answer by stanbon(75887)

Find the number of sides of a regular polygon if the measure of each interior angle is 17 times the measure of each exterior angle.
------
interior angle: 17x
exterior angle: x
----
Equation:
x + 17x = 180
18x = 180
x = 10 degrees
----
Sum of ALL the exterior angles is 360 degrees
----
# of exterior angles = 360/10 = 36
---
# of sides = # of exterior angles = 36
=======================================================
Cheers,
Stan H.
=============